# Algorithms with Calibrated Machine Learning Predictions

### v2

@@ 4.1 Setup
Prediction-aided scheduling.
@@
A prediction-aided algorithm $\mathcal{A}$ for job scheduling determines the probabilities $p_{i}$ from ML advice. Cho et al. (2022) assume access to a binary predictor $f_{b}:\mathcal{X}_{0}\to\{0,1\}$ of job priority and study the case where $p_{i}=\Pr[\vec{Y}_{i}=1\mid f_{b}(\vec{X}_{i})]$. These probabilities can be computed using Bayes’ rule, and because $f_{b}$ is binary, this procedure effectively assigns each job one of two probabilities. Although not explicitly discussed by Cho et al. (2022), this amounts to a basic form of post-hoc calibration. In contrast, our results extend to arbitrary calibrated predictors $f:\mathcal{X}_{0}\to[0,1]$—a more general framework that calls for new mathematical techniques—allowing us to significantly improve upon their results. In this setting, $\mathcal{A}$ takes the predictions $f(\vec{X})=\vec{v}$ as input and executes Algorithm 3 with probabilities $p_{i}=\vec{v}_{i}$.
@@
$\textnormal{ALG}(\mathcal{A},\vec{v},\vec{y})-\textnormal{OPT}(\vec{y})=\theta\omega_{1}L(\vec{v},\vec{y})+\theta\omega_{0}M(\vec{v},\vec{y})+(\omega_{1}-\omega_{0})N(\vec{v},\vec{y}),$ (3)
@@
Table 2: Quantities of interest in prediction-aided scheduling for fixed predictions $f(\vec{X})=\vec{v}$ and job priorities $\vec{y}$.
@@ 4.2 Scheduling with calibrated predictions
obtained by averaging the predictions of $f$ above and below the threshold $\beta$. Whereas $|R(f)|$ may be large, $f^{\prime}$ is only capable of outputting $|R(f^{\prime})|=2$ values. As a result, when ordering jobs with features $X_{1},\dots,X_{n}$ according to predictions from $f^{\prime}$, all jobs with $f(X)>\beta$ will be sequenced before jobs with $f(X)\leq\beta$, but the ordering of jobs within these bins will be random. In contrast, predictions from $f$ provide a more informative ordering of jobs (Figure 1). Note, however, that $f=f^{\prime}$ when $f$ has no variance in its predictions above or below the threshold $\beta$. We demonstrate in Theorem 4.3 that this intuition holds in general — improvements scale with the granularity of predictions.
@@
$\mathop{\mathbb{E}}\left[\mathbbm{1}_{\{\vec{Y}_{(i)}=0\}}\cdot\mathbbm{1}_{\{\vec{Y}_{(j)}=1\}}\mid f(\vec{X})\right]$ $=\biggl{(}\Pr[\vec{Y}_{(i)}=0\mid f(\vec{X}_{(i)})]\cdot\Pr[\vec{Y}_{(j)}=0\mid f(\vec{X}_{(j)})]\biggr{)}$
@@
$\mathop{\mathbb{E}}\left[\sum_{i=1}^{n}\sum_{j=i+1}^{n}(1-X_{(i)})\cdot X_{(j)}\right]\leq\mathop{\mathbb{E}}\left[\sum_{i=1}^{n}\sum_{j=i+1}^{n}(1-X_{i})\cdot X_{j}\right]$ $-\binom{n}{2}\cdot\mathrm{Var}(X_{1}).$
@@
With careful conditioning to deal with random summation bounds, we apply Lemma 4.2 to bound the expectations of $L(\cdot)$ and $N(\cdot)$, giving this section’s main theorem. Of note, Theorem 4.3 says that the expected number of inversions of high and low priority jobs decreases with predictor granularity, measured by $\kappa_{1}$ and $\kappa_{2}$. For the method from Cho et al. (2022), $\kappa_{1}=\kappa_{2}=0$ and the inequalities are tight.
@@
$\mathcal{A}$ is 1-consistent: $\textnormal{ALG}(\mathcal{A},\;\cdot)-\textnormal{OPT}(\cdot)=0$ when $\epsilon_{0}=\epsilon_{1}=0$.
@@ 5 Experiments
We now evaluate our algorithms on two real-world datasets, demonstrating the utility of using calibrated predictions. See Appendix C for additional details about our datasets and model training, as well a broader collection of results for different ML models and parameter settings.
@@ 5.1 Ski rental: Citi Bike rentals
To model the rent-or-buy scenario in the ski rental problem, we use publicly available Citi Bike usage data. [^1] Monthly usage data is publicly available at https://citibikenyc.com/system-data.. This dataset has been used for forecasting (Wang, 2016), system balancing (O’Mahony and Shmoys, 2015), and transportation policy (Lei and Ozbay, 2021), but to the best of our knowledge, this is its first use for ski rental. In this context, a Citi Bike user can choose one of two options: pay by ride duration (rent) or purchase a day pass (buy). If the user plans to ride for longer than the break-even point of $b$ minutes, it is cheaper to buy a day pass than to pay by trip duration. [^2] The day pass is designed to be more economical for multiple unlocks of a bike (e.g., $b\approx 66$ minutes for 1 unlock). However, ride data is anonymous, so we cannot track daily usage. We use single-ride durations to approximate the rent vs. buy trade-off for a spectrum of break-even points $b$. The distribution over ride durations can be seen in Appendix C.
@@
Figure 2: Comparison of $\mathop{\mathbb{E}}[\textnormal{ALG}/\textnormal{OPT}]$ for algorithms aided by predictions from a small MLP with two hidden layers of size 8 and 2. Algorithm 1 (Calibrated) performs best on average.
@@ 5.2 Scheduling: sepsis triage
We use a real-world dataset for sepsis prediction to validate our theory results for scheduling with calibrated predictions. Sepsis is a life-threatening response to infection that typically appears after hospital admission (Singer et al., 2016). Many works have studied using machine learning to predict the onset of sepsis, as every hour of delayed treatment is associated with a 4-8% increase in mortality (Kumar et al., 2006, Reyna et al., 2020); existing works aim to better predict sepsis to treat high-priority patients earlier. Replicating results from Chicco and Jurman (2020) we train a binary predictor for sepsis onset using logistic regression on a dataset of 110,204 hospital admissions. The base predictor achieves an AUC of 0.86 using age, sex, and septic episodes as features. We then calibrate this predictor using both the naive method from Cho et al. (2022) (Binary) and more nuanced histogram calibration (Zadrozny and Elkan, 2001) (Calibrated). Figure 3 shows the expected competitive ratio (normalized by the number of jobs $n=100$) achieved by Algorithm 3 when provided advice from each of these predictors for varying delay costs $\omega_{1},\omega_{0}$ and information barrier $\theta$. We see that the more nuanced predictions consistently result in schedules with smaller delay costs.
@@ 6 Conclusion
In this paper, we demonstrated that calibration is a powerful tool for algorithms with predictions, bridging the gap between traditional theoretical approaches—which rely on global uncertainty estimates—and modern ML methodologies that offer fine-grained, instance-specific uncertainty quantification. We focused on the ski rental and online scheduling problems, developing online algorithms that exploit local calibration guarantees to achieve strong average-case performance. For both problems, we highlighted settings where our algorithms outperform existing approaches, and we supported these theoretical findings with empirical evidence on real-world datasets. Beyond these two case studies, we believe calibration-based approaches offer broad potential for designing online decision-making algorithms, particularly in scenarios that require balancing worst-case robustness with reliable per-instance predictions.
Acknowledgements
Judy Hanwen Shen is supported by the Simons Foundation Collaboration on the Theory of Algorithmic Fairness, Anders Wikum is supported by a National Defense Science & Engineering Graduate (NDSEG) fellowship, and Ellen Vitercik acknowledges the support of NSF grant CCF-2338226. We thank Bailey Flanigan for stimulating early discussions that inspired us to pursue this research direction, and Ziv Scully for a technical insight in the proof of Lemma 4.2.
@@ Appendix A Ski Rental Proofs
$\leq 1+2\alpha+\min\left\{\mathop{\mathbb{E}}\left[f(X)\right]+\alpha,2\sqrt{\eta+K_{1}(f,\mathcal{D})+2\alpha}\right\}$ (Lemma 3.5)
$\leq 1+2\alpha+\min\left\{\mathop{\mathbb{E}}\left[f(X)\right]+\alpha,2\sqrt{\eta+3\alpha}\right\}.$
@@ Appendix B Scheduling Proofs
Given $n$ jobs to schedule with features $\vec{X}=(X_{1},\dots,X_{n})$ and the predictions $f(\vec{X})=(f(X_{1}),\dots,f(X_{n}))$, let $n_{1}=|\{i:f(X_{i})>\beta\}|$ be a random variable that counts the number of samples from $f$ with prediction larger than $\beta$. We’ll begin by computing expectations conditioned on $n_{1}$ before taking an outer expectation.
$\mathop{\mathbb{E}}[$ $L(f(\vec{X}),\vec{Y})\mid n_{1}]$
$=\mathop{\mathbb{E}}\biggl{[}\mathop{\mathbb{E}}[L(f(\vec{X}),\vec{Y})\mid f(\vec{X})]\mid n_{1}\biggr{]}$ (Tower property)
@@
$=\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=i+1}^{n_{1}}(1-f(\vec{X}_{(i)}))\cdot f(\vec{X}_{(j)})\mid n_{1}\right].$ (Calibration)
Performing the same computation for counts $M(\cdot)$ and $N(\cdot)$ gives
$\mathop{\mathbb{E}}[M(f(\vec{X}),\vec{Y})\mid n_{1}]=\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=i+1}^{n_{1}}(1-f(\vec{X}_{(i)}))\cdot(1-f(\vec{X}_{(j)}))\mid n_{1}\right]$
@@
$\mathop{\mathbb{E}}[N(f(\vec{X}),\vec{Y})\mid n_{1}]=$ $\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=n_{1}+1}^{n}(1-f(\vec{X}_{(i)}))\cdot f(\vec{X}_{(j)})\mid n_{1}\right]+$
$\mathop{\mathbb{E}}\left[\sum_{i=n_{1}+1}^{n}\sum_{j=i+1}^{n}(1-f(\vec{X}_{(i)}))\cdot f(\vec{X}_{(j)})\mid n_{1}\right].$
@@
$\mathop{\mathbb{E}}\biggl{[}$ $\sum_{i=1}^{n_{1}}\sum_{j=i+1}^{n_{1}}(1-f(\vec{X}_{(i)}))\cdot(1-f(\vec{X}_{(j)}))\mid n_{1}\biggr{]}$
$=\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=i+1}^{n_{1}}(1-f(\vec{X}_{i}))\cdot(1-f(\vec{X}_{j}))\mid n_{1}\right]$ (Lemma 4.1)
$=\binom{n_{1}}{2}\cdot\mathop{\mathbb{E}}\bigl{[}(1-f(\vec{X}_{i}))(1-f(\vec{X}_{j}))\mid f(\vec{X}_{i})>\beta,f(\vec{X}_{j})>\beta\bigr{]}$ (Independence)
$=\binom{n_{1}}{2}\cdot\Pr[Y=0\mid f(X)>\beta]^{2}$ (Calibration)
@@
$\mathop{\mathbb{E}}\biggl{[}$ $\sum_{i=1}^{n_{1}}\sum_{j=i+1}^{n_{1}}(1-f(\vec{X}_{(i)}))\cdot f(\vec{X}_{(j)})\mid n_{1}\biggr{]}$
$\leq\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=i+1}^{n_{1}}(1-f(\vec{X}_{i}))\cdot f(\vec{X}_{j})\mid n_{1}\right]-\binom{n_{1}}{2}\operatorname{Var}(f(X)\mid f(X)>\beta)$ (Lemma 4.2)
$=\binom{n_{1}}{2}\cdot\biggl{(}\mathop{\mathbb{E}}[(1-f(\vec{X}_{i}))f(\vec{X}_{j})\mid f(\vec{X}_{i}),f(\vec{X}_{j})>\beta]-\operatorname{Var}(f(X)\mid f(X)>\beta)\biggr{)}$
$=\binom{n_{1}}{2}\cdot\biggl{(}\Pr[Y=0\mid f(X)>\beta]\cdot\Pr[Y=1\mid f(X)>\beta]-\operatorname{Var}(f(X)\mid f(X)>\beta)\biggr{)}$
$=\binom{n_{1}}{2}\cdot\biggl{(}\frac{\rho(1-\rho)(1-\epsilon_{1})\epsilon_{0}}{\Pr[f(X)>\beta]^{2}}-\operatorname{Var}(f(X)\mid f(X)>\beta)\biggr{)}.$
@@
$\mathop{\mathbb{E}}\biggl{[}$ $\sum_{i=n_{1}+1}^{n}\sum_{j=i+1}^{n}(1-f(\vec{X}_{(i)}))\cdot f(\vec{X}_{(j)})\mid n_{1}\biggr{]}$
$\leq\mathop{\mathbb{E}}\left[\sum_{i=n_{1}+1}^{n}\sum_{j=i+1}^{n}(1-f(\vec{X}_{i}))\cdot f(\vec{X}_{j})\mid n_{1}\right]-\binom{n-n_{1}}{2}\operatorname{Var}(f(X)\mid f(X)\leq\beta)$
$=\binom{n-n_{1}}{2}\cdot\biggl{(}\Pr[Y=0\mid f(X)\leq\beta]\cdot\Pr[Y=1\mid f(X)\leq\beta]-\operatorname{Var}(f(X)\mid f(X)\leq\beta)\biggr{)}$
$=\binom{n-n_{1}}{2}\cdot\left(\frac{\rho(1-\rho)(1-\epsilon_{0})\epsilon_{1}}{\Pr[f(X)\leq\beta]^{2}}-\operatorname{Var}(f(X)\mid f(X)\leq\beta)\right).$
For the first term of $N(\cdot)$, we simply apply the rearrangement inequality in lieu of Lemma 4.2 for unordering, since the sum has the form $\sum_{i}a_{i}\cdot b_{i}$ where $a_{i}=(1-f(\vec{X}_{(i)}))$ is an increasing sequence and $b_{i}=\sum_{j=i+1}^{n}f(\vec{X}_{(j)})$ is a decreasing sequence.
$\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=n_{1}+1}^{n}(1-f(\vec{X}_{(i)}))\cdot f(\vec{X}_{(j)})\mid n_{1}\right]$ $\leq\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=n_{1}+1}^{n}(1-f(\vec{X}_{i}))\cdot f(\vec{X}_{j})\mid n_{1}\right]$
$=n_{1}(n-n_{1})\cdot\mathop{\mathbb{E}}[(1-f(\vec{X}_{i}))f(\vec{X}_{j})\mid f(\vec{X}_{i})>\beta,f(\vec{X}_{j})\leq\beta]$
@@
$\leq\mathop{\mathbb{E}}\left[\binom{n_{1}}{2}\cdot\biggl{(}\frac{\rho(1-\rho)(1-\epsilon_{1})\epsilon_{0}}{\Pr[f(X)>\beta]^{2}}-\operatorname{Var}(f(X)\mid f(X)>\beta)\biggr{)}\right]$
@@
$\leq\mathop{\mathbb{E}}\left[\binom{n-n_{1}}{2}\cdot\left(\frac{\rho(1-\rho)(1-\epsilon_{0})\epsilon_{1}}{\Pr[f(X)\leq\beta]^{2}}-\operatorname{Var}(f(X)\mid f(X)\leq\beta)\right)\right]$
@@
$\kappa_{1}:=\Pr[f(X)>\beta]^{2}\cdot\operatorname{Var}(f(X)\mid f(X)>\beta)\text{\quad and \quad}\kappa_{2}:=\Pr[f(X)\leq\beta]^{2}\cdot\operatorname{Var}(f(X)\mid f(X)\leq\beta).$

### v3

@@ 4.1 Setup
Learning-augmented scheduling.
@@
A learning-augmented algorithm $\mathcal{A}$ for job scheduling determines the probabilities $p_{i}$ from ML advice. Cho et al. (2022) assume access to a binary predictor $f_{b}:\mathcal{X}_{0}\to\{0,1\}$ of job priority and study the case where $p_{i}=\Pr[\vec{Y}_{i}=1\mid f_{b}(\vec{X}_{i})]$. These probabilities can be computed using Bayes’ rule, and because $f_{b}$ is binary, this procedure effectively assigns each job one of two probabilities. Although not explicitly discussed by Cho et al. (2022), this amounts to a basic form of post-hoc calibration. In contrast, our results extend to arbitrary calibrated predictors $f:\mathcal{X}_{0}\to[0,1]$—a more general framework that calls for new mathematical techniques—allowing us to significantly improve upon their results. In this setting, $\mathcal{A}$ takes the predictions $f(\vec{X})=\vec{v}$ as input and executes Algorithm 3 with probabilities $p_{i}=\vec{v}_{i}$.
@@
$\mathcal{A}(\vec{v},\vec{y})-\textnormal{OPT}(\vec{y})$ (3)
$=\theta\omega_{1}L(\vec{v},\vec{y})+\theta\omega_{0}M(\vec{v},\vec{y})+(\omega_{1}-\omega_{0})N(\vec{v},\vec{y}),$
@@
Table 2: Quantities of interest in learning-augmented scheduling for fixed predictions $f(\vec{X})=\vec{v}$ and job priorities $\vec{y}$.
@@ 4.2 Scheduling with calibrated predictions
obtained by averaging the predictions of $f$ above and below the threshold $\beta$. Whereas $|R(f)|$ may be large, $f^{\prime}$ is only capable of outputting $|R(f^{\prime})|=2$ values. As a result, when ordering jobs with features $X_{1},\dots,X_{n}$ according to predictions from $f^{\prime}$, all jobs with $f(X)>\beta$ will be sequenced before jobs with $f(X)\leq\beta$, but the ordering of jobs within these bins will be random. In contrast, predictions from $f$ provide a more informative ordering of jobs (Figure 1). Note, however, that $f=f^{\prime}$ when $f$ has no variance in its predictions above or below the threshold $\beta$. We demonstrate in Theorem 4.3 that this intuition holds in general: improvements scale with the granularity of predictions.
@@
$\mathop{\mathbb{E}}\left[\mathbbm{1}_{\{\vec{Y}_{(i)}=0\}}\cdot\mathbbm{1}_{\{\vec{Y}_{(j)}=1\}}\mid f(\vec{X})\right]$
$=\biggl{(}\Pr[\vec{Y}_{(i)}=0\mid f(\vec{X}_{(i)})]\cdot\Pr[\vec{Y}_{(j)}=0\mid f(\vec{X}_{(j)})]\biggr{)}$
@@
$\mathop{\mathbb{E}}\left[\sum_{i=1}^{n}\sum_{j=i+1}^{n}(1-X_{(i)})\cdot X_{(j)}\right]\leq$
$\mathop{\mathbb{E}}\left[\sum_{i=1}^{n}\sum_{j=i+1}^{n}(1-X_{i})\cdot X_{j}\right]$ $-\binom{n}{2}\cdot\mathrm{Var}(X_{1}).$
@@
With careful conditioning to deal with random summation bounds, we apply Lemma 4.2 to bound the expectations of $L(\cdot)$ and $N(\cdot)$, giving this section’s main theorem. Of note, Theorem 4.3 says that the expected number of inversions of high and low priority jobs decreases with predictor granularity, measured by $\kappa_{1}$ and $\kappa_{2}$. For the method from Cho et al. (2022), $\kappa_{1}=\kappa_{2}=0$ and the inequalities hold with equality.
@@
$\mathcal{A}(f(\vec{X}),\cdot)-\textnormal{OPT}(\cdot)=0$ when $\epsilon_{0}=\epsilon_{1}=0$, and $\mathcal{A}$ inherits the robustness guarantees of Cho et al. (2022) when $\epsilon_{0}$ and $\epsilon_{1}$ are large.
An analogous result holds under the weaker assumption that $f$ monotonically calibrated. That is, the empirical frequencies $\Pr[Y=1\mid f(X)]$ are non-decreasing in the prediction $f(X)$. This property holds trivially for calibrated predictors, but zero calibration error is not required. In fact, many calibration approaches used in practice (e.g. Platt scaling (Platt et al., 1999) and isotonic regression (Zadrozny & Elkan, 2001)) produce a monotonically calibrated predictor with non-zero calibration error. See Appendix B for details.
@@ 5 Experiments
We now evaluate our algorithms on two real-world datasets, demonstrating the utility of using calibrated predictions. See Appendix C for additional details about our datasets and model training, as well a broader collection of results for different ML models and parameter settings. [^1] Code and data available here: https://github.com/heyyjudes/algs-cali-pred
@@ 5.1 Ski rental: Citi Bike rentals
Figure 2: Comparison of $\mathbb{E}[\textnormal{ALG}/\textnormal{OPT}]$ for algorithms aided by predictions from a small MLP with two hidden layers of size 8 and 2. Algorithm 1 (Calibrated) performs best on average.
To model the rent-or-buy scenario in the ski rental problem, we use publicly available Citi Bike usage data. [^2] Monthly usage data is publicly available at https://citibikenyc.com/system-data.. This dataset has been used for forecasting (Wang, 2016), system balancing (O’Mahony & Shmoys, 2015), and transportation policy (Lei & Ozbay, 2021), but to the best of our knowledge, this is its first use for ski rental. In this context, a Citi Bike user can choose one of two options: pay by ride duration (rent) or purchase a day pass (buy). If the user plans to ride for longer than the break-even point of $b$ minutes, it is cheaper to buy a day pass than to pay by trip duration. [^3] The day pass is designed to be more economical for multiple unlocks of a bike (e.g., $b\approx 66$ minutes for 1 unlock). However, ride data is anonymous, so we cannot track daily usage. We use single-ride durations to approximate the rent vs. buy trade-off for a spectrum of break-even points $b$. The distribution over ride durations can be seen in Appendix C.
@@ 5.2 Scheduling: sepsis triage
We use a real-world dataset for sepsis prediction to validate our theory results for scheduling with calibrated predictions. Sepsis is a life-threatening response to infection that typically appears after hospital admission (Singer et al., 2016). Many works have studied using machine learning to predict the onset of sepsis, as every hour of delayed treatment is associated with a 4-8% increase in mortality (Kumar et al., 2006; Reyna et al., 2020); existing works aim to better predict sepsis to treat high-priority patients earlier. Replicating results from Chicco & Jurman (2020) we train a binary predictor for sepsis onset using logistic regression on a dataset of 110,204 hospital admissions. The base predictor achieves an AUC of 0.86 using age, sex, and septic episodes as features. We then calibrate this predictor using both the naive method from Cho et al. (2022) (Binary) and more nuanced histogram calibration (Zadrozny & Elkan, 2001) (Calibrated). Figure 3 shows the expected competitive ratio (normalized by the number of jobs $n=100$) achieved by Algorithm 3 when provided advice from each of these predictors for varying delay costs $\omega_{1},\omega_{0}$ and information barrier $\theta$. We see that the more nuanced predictions consistently result in schedules with smaller delay costs.
@@ 6 Conclusion
In this paper, we demonstrated that calibration is a powerful tool for algorithms with predictions in settings where performance is measured over a distribution and probabilistic estimates of a binary target enable good decisions. In particular, calibration bridges the gap between traditional theoretical approaches—which treat all predictions as equally reliable—and modern ML methodologies that offer fine-grained, instance-specific uncertainty quantification. We focused on the ski rental and online scheduling problems, developing online algorithms that exploit calibration guarantees to achieve strong average-case performance. For both problems, we highlighted settings where our algorithms outperform existing approaches and supported these findings with empirical evidence on real-world datasets.
This work exposes a number of directions for future research. For ski rental, deriving performance guarantees in terms of binary cross entropy and focusing on less rigid calibration measures (e.g. expected calibration error) offer to further close the gap between theory and practice. More broadly, we believe calibration-based approaches offer broad potential for designing online decision-making algorithms beyond these two case studies, particularly in scenarios that require balancing worst-case robustness with reliable per-instance predictions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgments
This work was supported in part by NSF grant CCF-2338226, the Simons Foundation Collaboration on the Theory of Algorithmic Fairness, and a National Defense Science & Engineering Graduate (NDSEG) fellowship. We thank Bailey Flanigan for stimulating early discussions that inspired us to pursue this research direction, and Ziv Scully for a technical insight in the proof of Lemma 4.2.
@@ Appendix A Ski Rental Proofs
$\leq 1+2\alpha+\min\left\{\mathop{\mathbb{E}}\left[f(X)\right]+\alpha,2\sqrt{\eta+3\alpha}\right\}.$ (Lemma 3.5)
@@ Appendix B Scheduling Proofs
We relax the calibration assumption and only assume that $f$ is monotonically calibrated, a weaker condition that the empirical frequencies $Z:=\Pr[Y=1\mid f(X)]$ are non-decreasing in the prediction $f(X)$. Given $n$ jobs to schedule with features $\vec{X}=(X_{1},\dots,X_{n})$ and the predictions $f(\vec{X})=(f(X_{1}),\dots,f(X_{n}))$, let $n_{1}=|\{i:f(X_{i})>\beta\}|$ be a random variable that counts the number of samples from $f$ with prediction larger than $\beta$, and define random variables $Z_{i}=\Pr[Y_{i}=1\mid f(X_{i})]$ which give empirical frequencies. We’ll begin by computing expectations conditioned on $n_{1}$ before taking an outer expectation.
$\mathop{\mathbb{E}}[L(f(\vec{X}),\vec{Y})\mid n_{1}]$ $=\mathop{\mathbb{E}}\biggl{[}\mathop{\mathbb{E}}[L(f(\vec{X}),\vec{Y})\mid f(\vec{X})]\mid n_{1}\biggr{]}$ (Tower property)
@@
$=\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=i+1}^{n_{1}}(1-Z_{(i)})Z_{(j)}\mid n_{1}\right].$
Performing the same computation for counts $M(\cdot)$ and $N(\cdot)$ yields
$\mathop{\mathbb{E}}[M(f(\vec{X}),\vec{Y})\mid n_{1}]=\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=i+1}^{n_{1}}(1-Z_{(i)})\cdot(1-Z_{(j)})\mid n_{1}\right]$
@@
$\mathop{\mathbb{E}}[N(f(\vec{X}),\vec{Y})\mid n_{1}]=\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=n_{1}+1}^{n}(1-Z_{(i)})\cdot Z_{(j)}\mid n_{1}\right]+\mathop{\mathbb{E}}\left[\sum_{i=n_{1}+1}^{n}\sum_{j=i+1}^{n}(1-Z_{(i)})\cdot Z_{(j)}\mid n_{1}\right].$
@@
$\mathop{\mathbb{E}}\biggl{[}$ $\sum_{i=1}^{n_{1}}\sum_{j=i+1}^{n_{1}}(1-Z_{(i)})\cdot(1-Z_{(j)})\mid n_{1}\biggr{]}$
$=\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=i+1}^{n_{1}}(1-Z_{i})\cdot(1-Z_{j})\mid n_{1}\right]$ (Lemma 4.1)
$=\binom{n_{1}}{2}\cdot\mathop{\mathbb{E}}\bigl{[}\Pr[Y=0\mid f(X)]\mid f(X)>\beta\bigr{]}^{2}$ (Independence)
$=\binom{n_{1}}{2}\cdot\Pr[Y=0\mid f(X)>\beta]^{2}$ (Tower property)
@@
$\mathop{\mathbb{E}}\biggl{[}$ $\sum_{i=1}^{n_{1}}\sum_{j=i+1}^{n_{1}}(1-Z_{(i)})\cdot Z_{(j)}\mid n_{1}\biggr{]}$
$\leq\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=i+1}^{n_{1}}(1-Z_{i})\cdot Z_{j}\mid n_{1}\right]-\binom{n_{1}}{2}\operatorname{Var}(Z\mid f(X)>\beta)$ (Lemma 4.2)
$=\binom{n_{1}}{2}\cdot\biggl{(}\Pr[Y=0\mid f(X)>\beta]\cdot\Pr[Y=1\mid f(X)>\beta]-\operatorname{Var}(Z\mid f(X)>\beta)\biggr{)}$
$=\binom{n_{1}}{2}\cdot\biggl{(}\frac{\rho(1-\rho)(1-\epsilon_{1})\epsilon_{0}}{\Pr[f(X)>\beta]^{2}}-\operatorname{Var}(Z\mid f(X)>\beta)\biggr{)}.$
@@
$\mathop{\mathbb{E}}\biggl{[}$ $\sum_{i=n_{1}+1}^{n}\sum_{j=i+1}^{n}(1-Z_{(i)})\cdot Z_{(j)}\mid n_{1}\biggr{]}$
$\leq\mathop{\mathbb{E}}\left[\sum_{i=n_{1}+1}^{n}\sum_{j=i+1}^{n}(1-Z_{i})\cdot Z_{j}\mid n_{1}\right]-\binom{n-n_{1}}{2}\operatorname{Var}(Z\mid f(X)\leq\beta)$ (Lemma 4.2)
$=\binom{n-n_{1}}{2}\cdot\biggl{(}\Pr[Y=0\mid f(X)\leq\beta]\cdot\Pr[Y=1\mid f(X)\leq\beta]-\operatorname{Var}(Z\mid f(X)\leq\beta)\biggr{)}$
$=\binom{n-n_{1}}{2}\cdot\left(\frac{\rho(1-\rho)(1-\epsilon_{0})\epsilon_{1}}{\Pr[f(X)\leq\beta]^{2}}-\operatorname{Var}(Z\mid f(X)\leq\beta)\right).$
For the first term of $N(\cdot)$, we simply apply the rearrangement inequality in lieu of Lemma 4.2 for unordering. Note that the sum has the form $\sum_{i}a_{i}\cdot b_{i}$, where $a_{i}=(1-Z_{(i)})$ and $b_{i}=\sum_{j=i+1}^{n}Z_{(j)}$. The sequence $\{a_{i}\}_{i=1}^{n}$ is non-decreasing as a result of the monotonic calibration of $f$, and $\{b_{i}\}_{i=1}^{n}$ is non-increasing. Thus,
$\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=n_{1}+1}^{n}(1-Z_{(i)})\cdot Z_{(j)}\mid n_{1}\right]$ $\leq\mathop{\mathbb{E}}\left[\sum_{i=1}^{n_{1}}\sum_{j=n_{1}+1}^{n}(1-Z_{i})\cdot Z_{j}\mid n_{1}\right]$
@@
$\leq\mathop{\mathbb{E}}\left[\binom{n_{1}}{2}\cdot\biggl{(}\frac{\rho(1-\rho)(1-\epsilon_{1})\epsilon_{0}}{\Pr[f(X)>\beta]^{2}}-\operatorname{Var}(Z\mid f(X)>\beta)\biggr{)}\right]$
@@
$\leq\mathop{\mathbb{E}}\left[\binom{n-n_{1}}{2}\cdot\left(\frac{\rho(1-\rho)(1-\epsilon_{0})\epsilon_{1}}{\Pr[f(X)\leq\beta]^{2}}-\operatorname{Var}(Z\mid f(X)\leq\beta)\right)\right]$
@@
$\kappa_{1}:=\Pr[f(X)>\beta]^{2}\cdot\operatorname{Var}(Z\mid f(X)>\beta)\text{\quad and \quad}\kappa_{2}:=\Pr[f(X)\leq\beta]^{2}\cdot\operatorname{Var}(Z\mid f(X)\leq\beta).$
The observation that $Z=f(X)$ when $f$ is calibrated gives the result from the main body. ∎